micha made a scale model of the empire state building. The building has an actual height of 381 meters. michas model used a scale in which 1 cm represents 50 meters. what is the height in centimeters of michas model?
Answer by addingup(3677)

1cm = 50
2 cm = 50 x 2
3cm = 50 x 3
and the inverse:
50m = 50/50 = 1cm
100m = 100/50 = 2cm
150m = 150/50 = 3cm
381m = 381/
you finish it.